Status Report

ISS On-Orbit Status 6 March 2002

By SpaceRef Editor
March 6, 2002
Filed under , ,

All ISS systems continue to function nominally, except as noted previously or below.

The crew received appreciative words for “slugging through the sound level test” yesterday and for their troubleshooting of the SSRMS.

More acoustic dosimetric measurements were taken today by FE-2 Dan Bursch [first collecting data from the yesterday’s survey, then deploying three static dosimeters in the Lab (one inside the TeSS with door closed), taking readings from them and then leaving them unattended for the sleep period until tomorrow. After again collecting readings, the equipment will be stowed].

During yesterday’s initial attempt to move the SSRMS/Canadarm2 out of the Lab window view for EARTHKAM (see below), the arm exhibited a failure signature last observed during Stage 6A. Preliminary analysis indicates that the SSRMS’s Wrist Roll Joint brake circuit may have a short, which prevents all SSRMS brakes from being disengaged on the primary electronics string and causes the arm to safe itself. The redundant (backup) string is unaffected. Analysis continues, but for tomorrow’s planned 8A S0 truss transfer dry run SSRMS operations would be run only on the redundant string. A failure on this string is not expected, but as a cautionary measure the SSRMS would no longer maneuver to positions which could interfere with Shuttle docking. For example, the dry run will not begin at the S0 grapple position in the Orbiter PLB (payload bay) but at the S0 PLB hover position. A review of tomorrow’s dry run was completed today, and the DOUG graphics display may have to be modified to account for the changes. An additional unrelated problem was observed on the RWS (robotics workstation) during the testing, which may be related to new software received with the R2 upgrade.

The ground began testing the recently installed R2 navigation software, and as of now the tests have been successful. They track and propagate Russian state vectors (position/velocity data), and the software will be active after the GPS receivers are installed on the ISS during the upcoming 8A mission.

CDR Yuri Onufrienko performed major (2.5 hr) maintenance on the SM’s condensate water processor (SRVK-2M), replacing the gas-liquid mixture filter (FGS) and filtration column assembly (BK) in the BKV water conditioning unit.

Yuri also set up the LIV video system, with lights, to record the activation of the ESA-sponsored GTS 1 (global timing system) payload, documenting the switch-on of its electronics unit for the ground (and customer) on a video cassette in the Russian Betacam VTR.

FE-1 Carl Walz continued work on the ARIS (active rack isolation system), which today returned the rack to a nominal configuration to resume operations. [Activities over the past few days focused first on those tasks that had to be done prior to the rotation of the ER2 rack, followed by the major rack rotation with actuator and pushrod maintenance, and concluding with adjustment of the snubber pins/cups and recalibration (which adjusted the upper actuators for proper pushrod positioning), ICE POP (isolation characterization experiment payload operations processor) reconfiguration, and finally alignment guide removal].

Bursch had been scheduled to prepare the EARTHKAM payload for preliminary operation, but the activity was deferred. Ê[The original plan of installing it again in the SM at window #8 had to be changed by MCC-H when Moscow voiced concern that there is no room for the system at that location, since the two pilot sights for testing/correcting/updating the BINS (Russian strap-down inertial guidance system) are currently installed at windows #8 and #6. EARTHKAM thus needs to return to Lab nadir window. Meanwhile, POIC (Payload Operations Integration Center) and MCC-H are working on the proper configuration to resolve the power issues of last January which had caused the relocation to the SM in the first place.]

Dan also performed work on the ADVASC (Advanced Astroculture) experiment, including a condensate fluid exchange and nutrient liquid exchange, along with taking second samples of the gas and condensate fluid in the growth chamber. The exchange activities replaced the removed fluid samples with fresh fluid.

After the ground had reconfigured BGA (beta gimbal assembly) 4B from Autotrack to Directed (“parked”) position at 125 degrees (BGA 2B remaining in Autotrack), station attitude control was handed over to the Russian motion control system (MCS) and maneuvered last night, at 7:39 pm EST, from XPOP (x-axis perpendicular to orbit plane) to LVLH (local vertical/local horizontal). A small adjustment then brought the ISS to the proper attitude for the two-burn reboost performed by the Progress 6P thrusters, controlled from the SM MCS via US-21 matching units. The maneuvers took place at 10:37 pm (2.5 minutes) and 11:29 pm EST (6.5 min.). Mean orbit altitude of the ISS was raised by about 6 km (absolute numbers see below). ISS attitude now remains in LVLH, as solar beta angle has been decreasing. Another two-burn reboost maneuver is planned for 3/13.

The ground is studying “excessive” leakage in the VRS (vacuum resources system). Along with the VES (vacuum exhaust system), VRS provides payloads in the Lab with a direct line to the vacuum of space. The system comprises vacuum lines, pressure sensors, vent valves, rack isolation valves and QDs (quick disconnects). The leak is suspected either in the payload rack or the mated QD connection. Additional troubleshooting will be timelined on 3/8 (Friday), during which the crew will perform checks on the QD and rack.

Carl Walz reconfigured the U.S. air conditioning system CCAA (common cabin air assembly) by switching over from the primary starboard channel to the alternate port channel. He also power-cycled the VOA (volatile organic analyzer), i.e., turning it off and on, with checkout, and later power-cycled the TEPC (tissue equivalent proportional counter) radiation instrumentation.

The crewmembers completed the weekly data collection for the NTXN Interactions (Vzaimodejstviye) experiment, regularly due on Wednesday. The activity consists of filling out a questionnaire on mood, group dynamics and journal observations, stored and backed up in an encrypted file on the HRF (human research facility) computer for later downlinking.

The crew performed their daily physical exercise regimen and performed routine maintenance tasks.

Onufrienko conducted a voice conference with the ground on the IMS (inventory management system).

Walz completed the periodic inspection and maintenance of the RED (resistive exercise device) and its bolts.

Carl later participated in an amateur (ham) radio chat with students at Deep Creek Elementary School in Boring, OR, a rural community not far from Mt. Hood. The school takes pride in having been designated recently as an “Exceptional School” by the Oregon Dept. of Education.

Today’s target areas for CEO observations were E. Mediterranean Dust and Smog (high pressure and fair weather hold over the eastern Med. On this pass, crew was to look to the right of track over Turkey for aerosols and any remains of yesterday’s Saharan dust cloud), Tigris-Euphrates, Turkey (pass ran the length of the Tigris-Euphrates Valley. Of interest: Beginning to map the lakes and reservoirs of the headwaters in Turkey, and if time permitted, continue mapping to the delta on the Persian Gulf), Lake Nasser, Toshka Lakes; Egypt (crew had an excellent opportunity to map this area, especially the new, rapidly changing, man-made lakes to west of the Nile River), Somalia Coast (ISS track was perpendicular to the coast. Crew was asked to use the ESC [electronic still camera] and try for near-nadir views of the dune pattern within the occasional breaks in the vegetation along the coast), Gulf of St. Lawrence (crew was to use this break in the weather to concentrate on ice conditions near Anticosti Island and the western coast of the isle of Newfoundland), Industrialized Southeastern Africa (high pressure persists in South Africa favoring aerosol buildup. As ISS entered this target area from the NW, crew was to try for oblique views to the right of track over the Val and Orange River valley industrialized region and continue taking photos as they crossed the coast), Eastern United States (best pass of the day for this target with cold high pressure in place. Of interest: oblique or limb shots either side of track from shortly after Lake Michigan to off shore from Cape Hatteras).

U.S. and Russian Segment Status (as of 2:05 pm EST today):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elekt2O2 generator is On (32 Amps mode). Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed). U.S. CDRA CO2 scrubber is in Standby.
  • BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, #2 in Purify mode.
  • SM Working Compartment: Pressure (mmHg) — 757, temperature (deg C) — 26.3, ppO2 (mmHg) — 148.3, ppCO2 (mmHg) — 2.5.
  • SM Transfer Compartment: Pressure (mmHg) — 762, temperature (deg C) — 21.4; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • FGB Cabin: Pressure (mmHg) — 756, temperature (deg C) — 20.3; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Node: Pressure (mmHg) — 748.34, temperature (deg C) — 21.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab: Pressure (mmHg) — 750.57, temperature (deg C) — n/a, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a;
  • Joint Airlock (Equip. Lock): Pressure (mmHg) — 750.37, temperature (deg C) — n/a; shell heater temp (deg C) — 23.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1: Shell heater temp (deg C) — 22.2.
  • PMA-2: Shell heater temp (deg C) — 18.4.

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment (USOS) not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]).

Electrical Power Systems (EPS):

  • Beta Gimbal Assembly (BGA) 2B in Autotrack mode, BGA 4B in Directed position at 125 degrees.
  • SM batteries: battery #1 is cycling; all other batteries (7) in “partial charge” mode.
  • FGB battery #5 is offline; battery #3 is cycling; all other batteries (4) are in “partial charge” mode.
  • Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:

  • Air conditioner SKV-1 is Off (Freon leak). SKV-2 is On.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is back-up, and C&C-3 is in standby, all with new R2 s/w.
  • GNC-2 MDM is prime; GNC-1 is back-up, with R1 s/w.
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Communications Systems:

  • S-band is operating nominally.
  • Ku-band is operating nominally in open loop pointing mode.
  • Audio subsystem operating nominally.
  • Video subsystem operating nominally.
  • MCOR (medium-rate communications outage recorder) operating nominally.


  • SSRMS/Canadarm2 at Port Stow position, with Keep Alive power on both strings.
  • RWS (robotics workstations) are Off.

ISS Orbit (as of this morning, 7:50am EST)

  • Mean altitude — 387.7 km
  • Apogee — 390.6 km
  • Perigee — 384.7 km
  • Period — 92.3 min.
  • Inclination (to Equator) — 51.64 deg
  • Eccentricity — 0.0004424
  • Orbits per 24-hr. day — 15.60
  • Altitude increase — 6000 m (mean) in last 24 hours
  • Solar Beta Angle: -37.0 deg (magnitude decreasing)
  • Revolutions since FGB/Zarya launch (Nov. ’98) — 18802
  • Current Flight Attitude — LVLH (local vertical/local horizontal [yaw: -10 deg, pitch: -7 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see

SpaceRef staff editor.